OpenStudy (anonymous):

limx->pi (-8x+sin7x)..... please help

OpenStudy (anonymous):

My answer was -0.4187909383 but for some reason the website keeps saying its wrong. Could you please help verify my answer....@TuringTest

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

maybe the website wants the exact value.... \(\huge \lim_{x \rightarrow \pi} (-8x+sin7x) = -8\pi + sin(7\pi)\) \(\huge =-8\pi + 0=-8\pi \)

OpenStudy (anonymous):

sorry, there is supposed to be a sin in front of the bracket....like this \[\lim_{x \rightarrow \pi}\sin(-8x+\sin7x)\]

OpenStudy (anonymous):

ok then... just take the sin(-8pi)....